Sasha owns two investments, A and B, that have a combined total value of 31,300 dollars. Investment A is expected to pay 24,300 dollars in 1 year(s) from today and has an expected return of 10.21 percent per year. Investment B is expected to pay 11,514 in 2 years from today and has an expected return of R per year. What is R, the expected annual return for investment B? Answer as a rate in decimal format so that 12.34% would be entered as .1234 and 0.98% would be entered as .0098.
Answer #1

FV in Investment A = 24,300
PV of Investment A = 24,300/(1+10.21%) = 22,048.81

PV of Investment B = 31,300 -22048.81 = 9251.1841

Rate of Return in B = (11,514/9251.1841)1/2-1 = 11.56% or 0.1156


Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well